OpenStudy (anonymous):

I'm at the moment studying complex numbers and having some troubles converting polar form to a+bi form. The problem is that I can't find the numerical representation of ex. sin 2pi/3 which is sqrt(3)/2. Is there any general way to figure this out or do I have to memorize the most common ones and hope I don't need more on future tests?

OpenStudy (anonymous):

@TuringTest

OpenStudy (turingtest):

you want a way to know that \(\sin\frac{2\pi}3=\frac{\sqrt3}2\) aside from memorization?

OpenStudy (anonymous):

Exactly

OpenStudy (turingtest):

there really isn't a good one, no. You just have to memorize the special values unfortunately...

OpenStudy (anonymous):

Ok, that was all i wanted to know, thank you! :)

OpenStudy (turingtest):

welcome!